2. Determine the convergence and divergence of the following series. (n + 3)! 3!n! 31
Answer #1

I am attaching the pictures of solution. If you find my answer helpful please press thumbs up button. Thanks

Ź (m+3)! E am = 31 m) 30 an= (m+3)! 31m 30 Now anti= (n+1+3)! 3! (MU! 3+) (replacem byn+i) - Anti 9 (m+4)! 31 (M+)3.3 Ratio tNow anti and an Х it Lol, then series is divergent. ③ il L=1, test is inconclusive. (m+4)! 31 (m+1)! 3 3 (using 6 (m+3)) 31anti Am (here (m+y) 3 (mt) may so for all nalo for all na >o for alln nu 1 (2+4) >o for all 3 (nt)) 1 (mt4) Imty) 3 (mt)) aslim (1+y) می (h n- A 30 (+1) - lim (1+) 3(+ (1+y) 3/1+ do) (1+0) 3(1to) if ktoj L - I 31+ Task As L= 1 / 3 = I () =) LCI 3.